# Gradient Descent for Noisy Optimization

### v1

@@ 5 Concluding remarks
The proposed algorithm (Algorithm 4) requires that GD-BLS can be called at most $J$ times, with $J$ a user-specified parameter. We conjecture that this condition can be removed and, assuming that A2 holds for $\alpha=\alpha^{\prime}$, that the resulting estimator $\hat{\theta}_{B}$ of the global minimiser $\theta_{\star}$ converges at speed $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-j_{B})}$ for some sequence $(j_{B})_{B\geq 1}$ such that $\lim_{B\rightarrow\infty}j_{B}=\infty$. Based on our numerical experiments, we conjecture that $j_{B}=a+b\log B$ for some constants $a\in[0,\infty)$ and $b\in(0,\infty)$ depending on the choice of the $\delta$ parameter of Algorithm 4. Confirming these results theoretically is left for future research.
Another interesting question that deserves further work is the possibility to replace, in Algorithm 4, GD-BLS by GD-BLS with Nesterov momentum (GD-BLS-NM). We recall the reader that, while the convergence of the former algorithm for computing $\min_{x\in\mathbb{R}^{d}}g(x)$ for a convex and $L$-smooth function $g$ is $1/T$, with $T$ the number of iterations, GD-BLS-NM converges to the global minimum of $g$ at the faster (and optimal) $1/T^{2}$ rate (see Nesterov,, 2018, Section 2.2). Based on this result, it is reasonable to expect that replacing GD-BLS by GD-BLS-NM will improve the non-asymptotic behaviour of the proposed procedure for computing $\theta_{\star}$. The extension of our convergence results to the case where GD-BLS-NM is used is however non-trivial, notably because it is unclear if the aforementioned fast convergence for this algorithm remains valid when the function $g$ is only locally $L$-smooth. Indeed, the key reason why the $1/T$ convergence rate of GD-BLS holds in this context is that the backtracking procedure ensures that the estimate of $\min_{x\in\mathbb{R}^{d}}g(x)$ improves from one iteration to the next, and this property of backtracking does not apply when Nesterov momentum is used.
@@ A.2.5 Proof of Lemma 4
A.2.6 Proof of Proposition 4
@@
A.2.7 Proof of Proposition 5
@@
then $v\geq\beta/L_{K}$. Since we trivially have $x_{t}\in\tilde{K}$ for all $t\geq 0$, this shows that $v_{t}\geq\beta/L_{K}$ for all $t\geq 1$ and the proof of the proposition is complete. ∎
A.2.8 Proof of Lemma 5
Proof of Lemma 5.
We now let
Using (26), we obtain that
and the result of the lemma follows.
@@ A.3.3 A second key property of Algorithm 3
To prove the result of the lemma let $L_{j+1,x}=\big{(}2\mathbb{E}[M^{\prime}_{\tilde{\Theta}_{j+1,x}}(Z)]\vee 1\big{)}$, $\theta_{0}\in\Theta_{j,x}$, $B\in\mathbb{N}$ and $\omega\in\Omega_{j,B}^{(2)}$. Remark that the function $F_{n_{B}}^{\omega}$ is strictly convex and $L_{j+1,x}$-smooth on the set $\tilde{\Theta}_{j+1,x}$, and that since $\omega\in\Omega_{j,B}^{(1)}$ we have $\theta^{\omega}_{\theta_{0},B,t}\in\Theta^{\prime}_{j+1,x}$ for all $t\in\{0,\dots,T_{\theta_{0},B}^{\omega}\}$. Next, assume first that $T_{\theta_{0},B}^{\omega}>0$ and for all $t\in\{0,\dots,T_{\theta_{0},B}^{\omega}-1\}$ let $v_{t}^{\omega}$ be such that $\theta^{\omega}_{\theta_{0},B,t+1}=\theta^{\omega}_{\theta_{0},B,t}+v_{t}^{\omega}\nabla F_{n_{B}}(\theta^{\omega}_{\theta_{0},B,t})$. Then, by using the computations done to proof Proposition 5 with $g=F_{n_{B}}^{\omega}$, $K=\tilde{\Theta}_{j+1,x}$, $\tilde{K}=\Theta^{\prime}_{j+1,x}$ and $L_{K}=L_{j+1,x}$, it follows that $v_{t}^{\omega}\geq\beta/L_{j+1}$ for all $t\in\{0,\dots,T_{\theta_{0},B}^{\omega}-1\}$ and thus that $C^{\omega}_{\theta_{0},B,t}\leq D_{j}n_{B}$, with $D_{j}=\beta^{k_{j}}+1$ where $k_{j}=\inf\{k\in\mathbb{N}:\,\beta^{k}<\beta/L_{j+1}\}$. In addition, by Proposition 4 we have

### v2

@@ 5 Concluding remarks
The proposed algorithm (Algorithm 4) requires that GD-BLS can be called at most $J$ times, with $J$ a user-specified parameter. We conjecture that this condition can be removed and, assuming that A2 holds for $\alpha=\alpha^{\prime}$, that the resulting estimator $\hat{\theta}_{B}$ of the global minimiser $\theta_{\star}$ converges at speed $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{j}_{B})}$ for some sequence $(j_{B})_{B\geq 1}$ such that $\lim_{B\rightarrow\infty}j_{B}=\infty$. Based on our numerical experiments, we predict that $j_{B}=a+b\log B$ for some constants $a\in[0,\infty)$ and $b\in(0,\infty)$ depending on the choice of $\delta$. Confirming these results theoretically is left for future research.
Another interesting question that deserves further work is the possibility to replace, in Algorithm 4, GD-BLS by GD-BLS with Nesterov momentum (GD-BLS-NM). We recall the reader that, while the convergence of the former algorithm for computing $\min_{x\in\mathbb{R}^{d}}g(x)$ for a convex and $L$-smooth function $g$ is $1/T$, with $T$ the number of iterations, GD-BLS-NM converges to the global minimum of $g$ at the faster (and optimal) $1/T^{2}$ rate (see Nesterov,, 2018, Section 2.2). Based on this result, it is reasonable to expect that replacing GD-BLS by GD-BLS-NM will improve the non-asymptotic behaviour of the proposed procedure for computing $\theta_{\star}$. The extension of our convergence results to the case where GD-BLS-NM is used is however non-trivial, notably because it is unclear if the aforementioned fast convergence rate of this algorithm remains valid when the function $g$ is only locally $L$-smooth. The $1/T$ convergence rate of GD-BLS holds for such functions because the backtracking procedure ensures that the estimate of $\min_{x\in\mathbb{R}^{d}}g(x)$ improves from one iteration to the next, and thus that the sequence in $\mathbb{R}^{d}$ generated by the algorithm stays in a compact set. By contrast, GD-BLS-NM is not a descent algorithm, and proving that the sequence in $\mathbb{R}^{d}$ it generates remains in compact set is non-trivial.
Finally, it is worth mentioning that, instead of backtracking line search, the procedure introduced by Malitsky and Mishchenko, (2019) can be used for choosing step-sizes $(v_{t})_{t\geq 1}$ which guarantee that the resulting gradient descent algorithm generates a sequence in $\mathbb{R}^{d}$ that stays in a compact set, when the objective function is convex. Under the additional assumption that the function $f(\cdot,Z)$ is $\mathbb{P}$-a.s. convex, it is easily checked that all the theoretical results presented in this work are valid if, instead of performing backtracking line search, in Algorithm 3 the step-sizes $(v_{t})_{t\geq 1}$ are chosen as proposed by Malitsky and Mishchenko, (2019). The key advantage of this approach is that, in the context of this work, it does not consume any computational budget and therefore allows to reduce the cost per iteration of Algorithm 3. However, this procedure has the drawback to require to choose the initial element $v_{1}$ of the sequence of $(v_{t})_{t\geq 1}$ which usually (and as confirmed by some unreported experiments) has a large impact on the behaviour of the algorithm.
@@ A.2.5 Proof of Lemma 4
A.2.6 Proof of Lemma 5
Proof of Lemma 5.
We now let
Using (21), we obtain that
and the result of the lemma follows.
A.2.7 Proof of Proposition 4
@@
A.2.8 Proof of Proposition 5
For a given $x_{0}\in\mathbb{R}^{d}$ the result of Proposition 5 follows by applying the following proposition with $\tilde{K}=\{x\in\mathbb{R}^{d}:\,g(x)\leq g(x_{0})\}$, noting that this set is indeed compact and convex since $g$ is assumed to be a convex function.
Proposition 6.
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex, differentiable and locally $L$-smooth function such that $g_{\star}>-\infty$. Let $x_{0}\in\mathbb{R}^{d}$, $\beta\in(0,1)$ and, for all $T\in\mathbb{N}$, let $x_{T}$ be as computed by Algorithm 1 where Algorithm 2 is used to compute $v_{t}$ for all $t\geq 1$. In addition, let $\tilde{K}\subset\mathbb{R}^{d}$ be a compact and convex set such that $x_{T}\in\tilde{K}$ for all $T\geq 1$,
and, noting that $K$ is a compact set, let $L_{K}\in[1,\infty)$ be such that $\|\nabla g(x)-\nabla g(x^{\prime})\|\leq L_{K}\|x-x^{\prime}\|$ for all $x,x^{\prime}\in K$. Then, (2) holds with $c=\beta/L_{K}$.
@@ A.2.8 Proof of Proposition 5
then $v\geq\beta/L_{K}$. Since by assumption $x_{T}\in\tilde{K}$ for all $T\geq 0$, this shows that $v_{T}\geq\beta/L_{K}$ for all $T\geq 1$ and the proof of the proposition is complete. ∎
@@ A.3.3 A second key property of Algorithm 3
To prove the result of the lemma let $L_{j+1,x}=\big{(}2\mathbb{E}[M^{\prime}_{\tilde{\Theta}_{j+1,x}}(Z)]\vee 1\big{)}$, $\theta_{0}\in\Theta_{j,x}$, $B\in\mathbb{N}$ and $\omega\in\Omega_{j,B}^{(2)}$. Remark that the function $F_{n_{B}}^{\omega}$ is strictly convex and $L_{j+1,x}$-smooth on the set $\tilde{\Theta}_{j+1,x}$, and that since $\omega\in\Omega_{j,B}^{(1)}$ we have $\theta^{\omega}_{\theta_{0},B,t}\in\Theta^{\prime}_{j+1,x}$ for all $t\in\{0,\dots,T_{\theta_{0},B}^{\omega}\}$.
Next, assume first that $T_{\theta_{0},B}^{\omega}>0$ and for all $t\in\{1,\dots,T_{\theta_{0},B}^{\omega}\}$ let $v_{t}^{\omega}$ be such that $\theta^{\omega}_{\theta_{0},B,t}=\theta^{\omega}_{\theta_{0},B,t-1}+v_{t}^{\omega}\nabla F_{n_{B}}(\theta^{\omega}_{\theta_{0},B,t-1})$. Then, by applying Proposition 6 with $g=F_{n_{B}}^{\omega}$, $K=\tilde{\Theta}_{j+1,x}$, $\tilde{K}=\Theta^{\prime}_{j+1,x}$ and $L_{K}=L_{j+1,x}$, it follows that $v_{t}^{\omega}\geq\beta/L_{j+1}$ for all $t\in\{1,\dots,T_{\theta_{0},B}^{\omega}\}$. From this result, we readily obtain that
$C^{\omega}_{\theta_{0},B,t}\leq D_{j}n_{B},\quad D_{j}=(\beta^{k_{j}}+1)(C_{\mathrm{eval}}+C_{\mathrm{grad}}),\quad k_{j}=\inf\{k\in\mathbb{N}:\,\beta^{k}<\beta/L_{j+1}\}.$
In addition, by Proposition 4 we have